# Rewriting Pre-Training Data Boosts LLM Performance in Math and Code

### v1

@@ 3.3.1 SGCR: Style-Guided Code Rewriting
SGCR enhances code readability by adding docstrings and type hints, unifying variable reassignment patterns, and standardizing function and class names in accordance with the Google Python Style Guide. Compared to the baseline (syntax-error and linter-based filtering), SGCR achieves improvements exceeding 9 points on both HumanEval and HumanEval+ as shown Figure 4. We also evaluate SGCR applied directly to the raw the-stack-v2-train-smol-ids corpus versus SGCR applied after syntax-error and linter-based filtering. As illustrated in Figure 4, the SGCR pipeline with syntax and linter filtering outperforms direct SGCR by approximately 2 points on downstream code-generation benchmarks. Consequently, we adopt the pipeline that syntax and linter-based filtering prior to SGCR in all subsequent experiments.
Ablation studies reveal that SGCR significantly improves HumanEval and HumanEval+ scores but results in an approximate 10-point decrease on the MBPP benchmark [9] (see Appendix G). Analysis indicates that MBPP’s canonical solutions often use non-standard function and class names, and SGCR’s automated renaming introduces identifier mismatches with MBPP’s test harness, leading to “is not defined” errors during evaluation. The identified mismatches obscure the model’s true code-generation capabilities, motivating us to exclude MBPP from our code-generation performance metrics across all experiments.
@@ 3.3.2 SCOR: Self-Contained Optimization Rewriting
While SGCR addresses stylistic inconsistencies, it does not modify program semantics. Manual inspection of models trained on SGCR-processed data reveals three recurring issues: (i) missing dependencies, where models attempt to import non-existent libraries or call undefined functions, causing runtime errors; (ii) inefficient algorithms, such as using naive recursive or quadratic approaches for tasks that support linear-time or dynamic-programming solutions; and (iii) trivial snippets, such as code that merely prints constants or performs basic arithmetic, offering minimal training value.
[…]
@@ 3.4 The Final SwallowCode dataset
Applying the complete pipeline—syntax-error filtering, Pylint-based filtering, Style-Guided Code Rewriting (SGCR), and Self-Contained Optimization Rewriting (SCOR)—to the-stack-v2-train-smol-ids produces the SwallowCode corpus, comprising 16.1 billion tokens. All intermediate artifacts, including non-optimal variants, are publicly available to support future research efforts.
@@
Figure 1 compares SwallowCode with several widely used open code datasets: CodeParrot-Clean (12.8 billion tokens) [^3] https://huggingface.co/datasets/codeparrot/codeparrot-clean, The Stack v1 (98.2 billion tokens) [1], The Stack v2-Smol (36.1 billion tokens) [2], and Stack-Edu (17.9 billion tokens) [3]. For a fair comparison, we extract only the Python subsets of each corpus and, following the protocol outlined in Section 3.1, allocate 8 billion code tokens within a 50 billion-token mixed batch, ensuring each sample is processed no more than once. SwallowCode outperforms all comparable publicly available corpora on the HumanEval (pass@1) and HumanEval+ (pass@1) benchmarks, demonstrating the effectiveness of our pipeline design. Detailed results are provided in Appendix H.1.
@@ 4 Construction of the math corpus
Section 3 demonstrated that LLM-driven rewriting significantly enhances the quality of code corpora. To evaluate the generalizability of this approach, we apply a tailored rewriting pipeline to mathematical data. We select finemath-4+ [3], a high-quality, publicly available math corpus, as the starting point and process it through a domain-specific rewriting pipeline. According to the evaluation in finemath [3], it outperforms other open mathematical datasets, including OpenWebMath [4], InfiMM-Math [28], and finemath-3+, in terms of quality on benchmarks such as GSM8K and MATH. Given its reported superior performance and public availability, we adopt finemath-4+ as the foundation corpus for constructing our mathematical corpus to maximize the effectiveness of our rewriting approach.
@@ 4.1 Experimental setup
We adhere to the protocol outlined in Section 3.1, performing continual pre-training of Llama-3.1-8B for approximately 50 billion tokens, varying only the target math corpus. The evaluation suite mirrors that of Section 3.1, with HumanEval+ replaced by the MATH dataset [29], comprising the following benchmarks: OpenBookQA, TriviaQA, HellaSwag, SQuAD 2.0, XWinO, MMLU, BBH, HumanEval, GSM8K, and MATH, with GSM8K and MATH as the primary math-focused benchmarks. The pre-training mixture comprises 82.2% multilingual text, 13.0% code, and 4.79% math; detailed proportions and data sources are provided in Appendix A.4.1. The complete hyperparameters are listed in the Appendix A.1.
@@ 4.2 LLM-driven rewriting
Figure 5: Performance gains from LLM-driven rewriting of finemath-4+ in a 50B-token continual pre-training ablation study. The rewritten corpus improves GSM8K(left) by 12.4 points and MATH(right) by 7.6 points.
The finemath-4+ corpus [3] contains a diverse range of data, from advanced mathematical problems to basic arithmetic, alongside extraneous non-mathematical documents. This heterogeneity renders rule-based filtering challenging, as it struggles to distinguish relevant mathematical content from irrelevant artifacts. To address this, we design an LLM-driven rewriting pipeline using Llama-3.3-70B-Instruct, which not only cleans and refines the data but also enhances its quality for mathematical reasoning tasks.
The pipeline instructs the model to: (1) remove residual web headers, footers, and privacy notices; (2) eliminate extraneous metadata, such as question and answer timestamps; (3) restore missing context in incomplete questions or answers; (4) rewrite explanations to be concise yet comprehensive; and (5) provide clear, step-by-step solutions. Steps (1) and (2) mirror the syntax error and linter-based filtering applied to SwallowCode (Section 3), addressing inappropriate content that rule-based methods alone could not effectively filter. Steps (3) through (5) parallel the self-containment and style enhancements of the code rewriting pipeline, adapting them to the mathematical domain. The complete prompt is provided in Appendix E.
As shown in Figure 5, the rewritten corpus yields substantial improvements: 12.4 points on GSM8K and 7.6 points on MATH in a 50 billion tokens continual pre-training ablation experiment. These results demonstrate that LLM-driven rewriting, while tailored to the unique characteristics of mathematical data, successfully enhances the already high-quality finemath-4+ corpus. This confirms the generalizability of our rewriting approach beyond code, offering a robust method for improving open-domain datasets for mathematical reasoning.
@@ 5 Limitations
While SwallowCode and SwallowMath significantly improve code generation and mathematical reasoning, several limitations should be noted. First, the rewriting pipelines may preserve biases present in the source data. For example, the-stack-v2-train-smol-ids may over-represent certain coding patterns (e.g., recursive algorithms over iterative ones), and Finemath-4+ may favor specific problem types (e.g., algebra over geometry). Additionally, as the rewriting process relies on Llama-3.3-70B-Instruct, the resulting datasets may reflect this model’s preferences, such as favoring certain variable naming conventions or solution strategies, potentially limiting generalizability. Second, our evaluation is confined to continual pre-training with a 50 billion token budget, as detailed in Section 3.1, to ensure controlled and reproducible experiments within computational constraints. The impact of extending pre-training beyond this budget remains unexplored, and performance trends at larger scales may differ, particularly for tasks requiring extensive training data. Third, although the SwallowCode pipeline is designed to be language-agnostic, requiring only static syntax checking and linter tools, our experiments focus exclusively on Python to facilitate automated evaluation. Empirical validation for other languages (for example, Java, C++) was not feasible due to resource constraints, limiting evidence of the pipeline’s broader applicability.
@@ 6 Conclusion
We introduce SwallowCode and SwallowMath, two openly licensed corpora under the Llama 3.3 Community License, designed to elevate large language model (LLM) performance in program synthesis and mathematical reasoning through systematic data rewriting. Unlike prior approaches that rely on exclusionary filtering (e.g., Stack-Edu [3]) or limited transformations (e.g., Jain et al. [12]), our transform-and-retain paradigm upgrades low-quality data, maximizing its utility. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 [2] using a four-stage pipeline: syntax validation, pylint-based style filtering, Style-Guided Code Rewriting (SGCR), and Self-Contained Optimization Rewriting (SCOR). SwallowMath ($\approx$2.3 billion tokens) enhances Finemath-4+ [3] by removing boilerplate, restoring context, and reformatting solutions into concise, step-by-step explanations.
Continual pre-training of Llama-3.1-8B [5] for 50 billion tokens with SwallowCode surpasses the baseline model’s code generation capabilities, achieving pass@1 gains of +17.0 on HumanEval and +17.7 on HumanEval+ compared to Stack-Edu. Similarly, substituting SwallowMath yields accuracy improvements of +12.4 on GSM8K and +7.6 on MATH. Ablation studies confirm that each pipeline stage contributes incrementally, with the rewriting stages providing the most significant boosts. Rigorous checks ensure no test-set leakage, reinforcing the robustness of these results.
While our experiments focus on Python for automated evaluation, the pipeline’s design is language-agnostic, requiring only parseable syntax and a linter, as discussed in Section 3.4. To enable reproducible research and foster advancements in LLM capabilities, we release all datasets, prompts, and checkpoints. SwallowCode and SwallowMath establish a scalable framework for high-quality data curation, paving the way for enhanced automated reasoning and software development.
@@ Appendix B Code linting filtering
When conducting linting with pylint, warnings and errors dependent on the linting environment, such as import errors or line length violations, are excluded using the --disable option. Additionally, some files with Python extensions primarily contain comments with textual content and have minimal script functionality. To address this, as discussed in Section 3.2.2, we introduced a mechanism that imposes a heuristic penalty based on the proportion of comments to filter out such files.
@@ F.1 Computational cost of LLM scoring
The data synthesis process utilized vLLM 0.7.2 and PyTorch 2.5.1. The dataset was synthesized with a global batch size of 2048 and tensor parallelism of 4. Data generation was performed using four H100 (94 GB) GPUs per job. With an input processing speed of approximately 2000 tokens/s and an output generation speed of approximately 3000 tokens/s, and given an average input dataset length of 836 tokens, an average output dataset length of 1271 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.2.3 consumed a total of 4,869 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
@@ F.2 Computational cost of LLM rewriting
The synthesis process for LLM rewriting also employed vLLM 0.7.2 and PyTorch 2.5.1, with a global batch size of 2048 and tensor parallelism of 4. Data generation was conducted using 4 H100 (94GB) GPUs per job. The input processing speed was approximately 2000 tokens/s, and the output generation speed was approximately 3000 tokens/s. Given an average input dataset length of 836 tokens, an average output dataset length of 1819 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.3.1 consumed a total of 5,925 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
@@ H.1 Code ablation experiments results
As described in Section 3.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. In the following, we report the evaluation results for 13 code ablation experiments, with their relationships illustrated in Figure 10. Experiments exp1, exp8, exp9, and exp13 serve as baselines, utilizing only Python data extracted from existing open code corpora. The remaining experiments are conducted to construct the SwallowCode dataset. We evaluated performance using the following nine benchmarks: OpenBookQA [17], TriviaQA [18], HellaSwag [19], SQuAD 2.0 [20], XWinograd [21], MMLU [22], GSM8K [23], BBH [24], HumanEval [25], and HumanEval+ [26].
@@ H.2 Math ablation experiments results
As described in Section 4.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. Below, we present the evaluation results for two code ablation experiments. Specifically, we adopted the following ten evaluation benchmarks: OpenBookQA [17], TriviaQA [18], HellaSwag [19], SQuAD 2.0 [20], XWinograd [21], MMLU [22], GSM8K [23], BBH [24], HumanEval [25], and MATH [22].

### v2

@@ 3.3.1 SGCR: Style-Guided Code Rewriting
SGCR enhances code readability by adding docstrings and type hints, unifying variable reassignment patterns, and standardizing function and class names in accordance with the Google Python Style Guide. Compared to the pre-rewriting (syntax-error and linter-based filtering), SGCR achieves improvements over 9 points on both HumanEval and HumanEval+ as shown Figure 4. We also evaluate SGCR applied directly to the raw the-stack-v2-train-smol-ids corpus versus SGCR applied after syntax-error and linter-based filtering. As illustrated in Figure 4, the SGCR pipeline with syntax and linter filtering outperforms direct SGCR by approximately 2 points on downstream code-generation benchmarks. Consequently, we adopt the pipeline that syntax and linter-based filtering prior to SGCR in all subsequent experiments.
Ablation studies reveal that SGCR significantly improves HumanEval and HumanEval+ scores but results in an approximate 10-point decrease on the MBPP benchmark [12] (detailed in Appendix G). Analysis indicates that MBPP’s solutions often use non-standard function and class names, and SGCR’s automated renaming introduces function name mismatches with MBPP’s unit tests, leading to “undefined” errors during evaluation. The identified mismatches obscure the model’s true code-generation capabilities, motivating us to exclude MBPP from our evaluation benchmarks across all experiments.
@@ 3.3.2 SCOR: Self-Contained Optimization Rewriting
Although SGCR ensures adherence to stylistic criteria, it does not modify the program semantics. Manual observation of models trained on SGCR-processed data reveals three recurring issues: (i) missing dependencies, where models attempt to import non-existent libraries or call undefined functions, causing runtime errors; (ii) inefficient algorithms, such as using naive recursion or quadratic-order algorithms for problems that linear-order or dynamic-programming solutions; and (iii) trivial snippets, such as code that merely prints constants or performs basic arithmetic, offering minimal training value.
To address these limitations while preserving SGCR’s stylistic improvements, we introduce Self-Contained Optimization Rewriting (SCOR). Guided by a ten-rule prompt (detailed in Appendix D.2), SCOR rewrites each snippet to ensure self-containment by inlining or satisfying external dependencies, replaces inefficient algorithms with more computationally efficient alternatives, and transforms trivial code into meaningful, executable examples. As illustrated in Figure 4, SCOR improves HumanEval and HumanEval+ scores by over 5 points compared to SGCR. These results underscore the importance of semantic-level rewrites beyond stylistic enhancements, establishing SCOR as the final stage of the SwallowCode construction pipeline. We did not conduct an ablation experiment evaluating SCOR in isolation without SGCR. However, prompt validation experiments with Llama-3.3-70B-Instruct indicated that simultaneously applying SGCR and SCOR often reduced the quality of the rewritten code due to the challenges LLMs face in balancing multiple objectives. This led to the adoption of a two-stage SGCR and SCOR approach, with the isolated SCOR evaluation deferred to future work.
Compared to the 2-point performance gain achieved during the filtering stages (green line in Figure 4), the rewriting stages using SGCR and SCOR led to a total performance improvement of 14 points—9 points from SGCR and 5 points from SCOR. This highlights the significant potential for enhancing dataset curation by incorporating LLM-driven rewriting approach.
@@ 3.4 The Final SwallowCode dataset
Applying the complete pipeline, including syntax error filtering, Pylint-based filtering, Style-Guided Code Rewriting (SGCR), and Self-Contained Optimization Rewriting (SCOR), to the-stack-v2-train-smol-ids produces the SwallowCode corpus, comprising 16.1 billion tokens. All intermediate artifacts, including non-optimal variants, are publicly available to support future research efforts.
@@
Figure 1 compares SwallowCode with several widely used open code datasets: CodeParrot-Clean (12.8 billion tokens) [^3] https://huggingface.co/datasets/codeparrot/codeparrot-clean, The Stack v1 (98.2 billion tokens) [1], The Stack v2-Smol (36.1 billion tokens) [2], and Stack-Edu (17.9 billion tokens) [3]. For a fair comparison, we extract only the Python subsets of each corpus and, following the protocol outlined in Section 3.1, allocate 16% (8 billion) code tokens within a 50 billion-token mixed batch, ensuring each sample is processed no more than once. SwallowCode outperforms all comparable publicly available corpora on the HumanEval (pass@1) and HumanEval+ (pass@1) benchmarks, demonstrating the effectiveness of our pipeline design. Detailed results are provided in Appendix H.1.
@@ 4 Construction of the math corpus
Section 3 demonstrated that LLM-driven rewriting significantly boosts coding performance. To evaluate the transferability of this approach, we apply a tailored rewriting pipeline to mathematical data. We select finemath-4+ [3], a high-quality, publicly available math corpus, as the starting point and process it through a rewriting pipeline. According to the evaluation in finemath [3], it outperforms other open mathematical datasets, including OpenWebMath [4], InfiMM-Math [33], and finemath-3+, in terms of the performance on benchmarks such as GSM8K and MATH. Given its reported superior performance and public availability, we adopt finemath-4+ as the foundation corpus for constructing our mathematical corpus to maximize the effectiveness of our rewriting approach.
@@ 4.1 Experimental setup
We adhere to the protocol outlined in Section 3.1, performing continual pre-training of Llama-3.1-8B for approximately 50 billion tokens, varying only the target math corpus. The evaluation benchmarks mirror those of Section 3.1, with HumanEval+ replaced by the MATH dataset [34], with GSM8K and MATH as the primary math-focused benchmarks. The pre-training mixture comprises 82.2% multilingual text, 13.0% code, and 4.79% math; detailed proportions and data sources are provided in Appendix A.4.2. The complete hyperparameters are listed in the Appendix A.1.
@@ 4.2 LLM-driven rewriting
Figure 5: Performance gains from LLM-driven rewriting of finemath-4+. The rewritten corpus improves GSM8K(left) by 12.4 points and MATH(right) by 7.6 points.
The Finemath‑4+ corpus [3] is a collection of documents in which snippets of mathematical text are embedded within passages that are otherwise unrelated to mathematics. In addition, the mathematical content ranges widely, from elementary arithmetic to advanced topics. This heterogeneity renders rule-based filtering challenging, as it struggles to distinguish relevant mathematical content from irrelevant artifacts. To address this, we design an LLM-driven rewriting pipeline using Llama-3.3-70B-Instruct, which not only cleans and refines the data but also enhances its quality for mathematical reasoning tasks.
The rewriting prompt instructs the model to: (1) remove residual web headers, footers, and privacy notices; (2) eliminate irrelevant metadata, such as question and answer timestamps; (3) restore missing context in incomplete questions or answers; (4) rewrite derivation steps to be concise yet comprehensive; and (5) provide clear, step-by-step solutions. Steps (1) and (2) are analogous to the syntax error and linter-based filtering applied to SwallowCode (Section 3), addressing inappropriate content that rule-based methods alone could not effectively filter. Steps (3) through (5) parallel the self-containment and style enhancements of the code rewriting pipeline, adapting them to the mathematical domain. The complete prompt is provided in Appendix E.
As shown in Figure 5, the rewritten corpus yields substantial improvements: 12.4 points on GSM8K and 7.6 points on MATH. These results demonstrate that LLM-driven rewriting, while tailored to the unique characteristics of mathematical data, successfully enhances the already high-quality finemath-4+ corpus. This confirms the generalizability of our rewriting approach beyond code, offering a robust method for improving open-domain datasets for mathematical reasoning.
@@ 5 Limitations
While SwallowCode and SwallowMath significantly improve code generation and mathematical reasoning performance of LLMs, several limitations should be noted. First, the rewriting pipelines may preserve biases present in the source data. For example, the-stack-v2-train-smol-ids may over-represent certain coding patterns (e.g., recursive algorithms over iterative ones), and Finemath-4+ may favor specific problem types (e.g., algebra over geometry). Additionally, as the rewriting process relies on Llama-3.3-70B-Instruct, the resulting datasets may reflect this model’s preferences, such as favoring certain variable naming conventions or solution strategies, potentially limiting generalizability. Second, our evaluations are confined to continual pre-training with a 50 billion token budget, as detailed in Section 3.1, to ensure controlled and reproducible experiments within computational constraints. The impact of extending pre-training beyond this budget remains unexplored, and performance trends at larger scales may differ, particularly for tasks requiring extensive training data. Third, although the SwallowCode pipeline is designed to be language-agnostic, requiring only static syntax checking and linter tools, our experiments focus exclusively on Python to facilitate automated evaluation. Empirical validation for other languages (for example, Java, C++) was not feasible due to resource constraints, limiting evidence of the pipeline’s broader applicability.
@@ 6 Conclusion
In this work, we present SwallowCode and SwallowMath, two pre-training datasets released under the Llama 3.3 Community License, designed to enhance large language model (LLM) capabilities in program synthesis and mathematical reasoning. SwallowCode, spanning approximately 16.1 billion tokens, refines Python code from The-Stack-v2 using a transform-and-retain pipeline that includes syntax validation, style-guided rewriting, and self-contained optimization. SwallowMath, with approximately 2.3 billion tokens, upgrades Finemath-4+ by reformatting mathematical content into concise, context-rich solutions. Unlike traditional approaches that discard low-quality data, our method systematically enhances it, maximizing utility and setting a new standard for dataset curation.
SwallowCode and SwallowMath achieve state-of-the-art performance among publicly available datasets for code and math reasoning, demonstrating the efficacy of our data refinement strategies. The language-agnostic design of our pipeline, which requires only parseable syntax and a linter, suggests its potential applicability to other programming languages. To foster reproducible research and drive advancements in LLM pretraining, we release SwallowCode, SwallowMath, all associated prompts, and model checkpoints. These resources establish a scalable framework for high-quality data curation, paving the way for enhanced automated reasoning and software development.
@@ A.4 Data mixture for ablation experiments
The SwallowProject [^6] https://swallow-llm.github.io/index.en.html, a collaborative effort between the Institute of Science Tokyo, Japan, and the National Institute of Advanced Industrial Science and Technology (AIST), aims to develop open-source large language models (LLMs) with strong capabilities in both Japanese and English. A key challenge in continual pre-training from high-performing models like Llama-3.1-8B is mitigating catastrophic forgetting, particularly in maintaining or improving mathematical reasoning and code generation performance. To address this, our ablation experiments were designed to enhance Llama-3.1-8B’s performance on HumanEval, HumanEval+, GSM8K, and MATH while incorporating a multilingual dataset predominantly composed of Japanese and English text. This approach led to the development of SwallowCode and SwallowMath, as detailed in Sections 3 and 4. The data mixture reflects a high proportion of Japanese text, consistent with our project’s focus on bilingual proficiency. However, as described in Sections 3.1 and 4.1, all ablation experiments maintain identical settings except for the target code or math dataset, ensuring a fully controlled experimental design.
This continual pre-training strategy aligns with established practices in high-quality LLM development, as seen in models like OLMo-2 [5] and Nemotron-H [6]. Specifically, adopting a staged pre-training approach, as observed in OLMo-2 and Nemotron-H, where later training phases leverage high-quality multilingual text alongside specialized math and code datasets to enhance LLM capabilities, ensures that our ablation experiments align with realistic LLM training scenarios.
@@ Appendix B Code linting filtering
When conducting linting with pylint, warnings and errors dependent on the linting environment, such as import errors, are excluded using the --disable option. Additionally, some files with Python extensions primarily contain comments with textual content and have minimal script functionality. To address this, as discussed in Section 3.2.2, we introduced a mechanism that imposes a heuristic penalty based on the proportion of comments to filter out such files.
@@ F.1 Computational cost of LLM scoring
The data synthesis process utilized vLLM 0.7.2 and PyTorch 2.5.1. The dataset was synthesized with a global batch size of 2048 and tensor parallelism of 4. Data generation was performed using four H100 (94 GB) GPUs per job. With an input processing speed of approximately 2000 tokens/s and an output generation speed of approximately 3000 tokens/s, and given an average input dataset length of 836 tokens, an average output dataset length of 1271 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.2.3 consumed a total of 19,477 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
@@ F.2 Computational cost of LLM rewriting
The synthesis process for LLM rewriting also employed vLLM 0.7.2 and PyTorch 2.5.1, with a global batch size of 2048 and tensor parallelism of 4. Data generation was conducted using 4 H100 (94GB) GPUs per job. The input processing speed was approximately 2000 tokens/s, and the output generation speed was approximately 3000 tokens/s. Given an average input dataset length of 836 tokens, an average output dataset length of 1819 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.3.1 consumed a total of 23,703 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
@@ H.1 Code ablation experiments results
As described in Section 3.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. In the following, we report the evaluation results for 13 code ablation experiments, with their relationships illustrated in Figure 10. Experiments exp1, exp8, exp9, and exp13 serve as baselines, utilizing only Python data extracted from existing open code corpora. The remaining experiments are conducted to construct the SwallowCode dataset. We evaluated performance using the following ten benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWinograd [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and HumanEval+ [31].
@@ H.2 Math ablation experiments results
As described in Section 4.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. Below, we present the evaluation results for two math ablation experiments. Specifically, we adopted the following ten evaluation benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWinograd [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and MATH [27].